OpenStudy (anonymous):

Find the perimeter of a square having two consecutive vertices at (7, 2) and (-3, 3).

jimthompson5910 (jim_thompson5910):

Find the distance between those two points

jimthompson5910 (jim_thompson5910):

That will give you the length of each side

jimthompson5910 (jim_thompson5910):

Once you know the length of each side, you multiply by 4 to get the perimeter

OpenStudy (anonymous):

so how do I exactly find the length im horrible at math.

jimthompson5910 (jim_thompson5910):

you use the distance formula d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 )

jimthompson5910 (jim_thompson5910):

the first point is (x1,y1) in our case, the first point is (7,2)

jimthompson5910 (jim_thompson5910):

so (x1,y1) = (7,2)

jimthompson5910 (jim_thompson5910):

which means x1 = 7 y1 = 2

jimthompson5910 (jim_thompson5910):

similarly, the second point is (-3, 3) so x2 = -3 y2 = 3

jimthompson5910 (jim_thompson5910):

plug all that in and simplify d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) d = sqrt( (2-7)^2 + (3-(-3))^2 ) d = sqrt( (2-7)^2 + (3+3)^2 ) I'll let you finish

jimthompson5910 (jim_thompson5910):

oh wait, made a serious typo

jimthompson5910 (jim_thompson5910):

let me fix it

jimthompson5910 (jim_thompson5910):

d = sqrt( (x2 - x1)^2 + (y2 - y1)^2 ) d = sqrt( (-3 - 7)^2 + (3 - 2)^2 ) Ok fixed

OpenStudy (anonymous):

thank you so much

jimthompson5910 (jim_thompson5910):

tell me what you get

OpenStudy (anonymous):

101

jimthompson5910 (jim_thompson5910):

now take the square root of that to get the exact distance to be \[\Large d = \sqrt{101}\]

jimthompson5910 (jim_thompson5910):

so the exact perimeter would be \[\Large P = 4\sqrt{101}\] since you multiply the side length by 4 to get the perimeter

jimthompson5910 (jim_thompson5910):

you can use a calculator to get the approximate answer

OpenStudy (anonymous):

that's what the answer is suppose to be thank you so much

jimthompson5910 (jim_thompson5910):

yw